# Soundscape Captioning using Sound Affective Quality Network and Large Language Model

### v2

@@ Abstract
We live in a rich and varied acoustic world, which is experienced by individuals or communities as a soundscape. Computational auditory scene analysis, disentangling acoustic scenes by detecting and classifying events, focuses on objective attributes of sounds, such as their category and temporal characteristics, ignoring their effects on people, such as the emotions they evoke within a context. To fill this gap, we propose the soundscape captioning task, which enables automated soundscape analysis, thus avoiding labour-intensive subjective ratings and surveys in conventional methods. With soundscape captioning, context-aware descriptions are generated for soundscape by capturing the acoustic scene, event information, and the corresponding human affective qualities (AQs). To this end, we propose an automatic soundscape captioner (SoundSCaper) system composed of an acoustic model, i.e. SoundAQnet, and a large language model (LLM). SoundAQnet simultaneously models multi-scale information about acoustic scenes, events, and perceived AQs, while the LLM describes the soundscape with captions by parsing the information captured with SoundAQnet. The soundscape caption’s quality is assessed by a jury of 16 audio/soundscape experts. The average score (out of 5) of SoundSCaper-generated captions is lower than the score of captions generated by two soundscape experts by 0.21 and 0.25, respectively, on the evaluation set and the model-unknown mixed external dataset with varying lengths and acoustic properties, but the differences are not statistically significant. Overall, the proposed SoundSCaper shows promising performance, with captions generated being comparable to those annotated by soundscape experts. The code of models, LLM scripts, human assessment data and instructions, and expert evaluation statistics are all publicly available.
@@ I Introduction
The definition of soundscape in ISO 12913-1:2014 [1]: “the acoustic environment as perceived or experienced and/or understood by a person or people, in context”, emphasizes the interaction between the person and the acoustic environment. Recognition of salient individual audio events contributes to understanding and experiencing the meanings and associations they evoke, a primary cognitive process [2]. Individual sounds that stand out for their sensory salience contribute to perceptual dimensions such as pleasure [3]. Sensory salience, amongst others, depends on the loudness of the sound. The recognized sounds, together with the more general background, also trigger an emotional pathway that leads to an experience of pleasure and eventfulness [2]. This appraisal happens within the context of creating expectations. Violation of these expectations affects the detection of sound and might impact the appraisal of the sonic environment [4]. Context and situational awareness are formed via audio or visual cues and prior knowledge about the place, either from prior experience or associations with prototypes. Thus, recognizing the auditory scene also plays a vital role in forming perceptions and understandings of acoustic environments. In summary, AI models for soundscape analysis, including perceived acoustic quality, will benefit from the recognition of acoustic scenes and audio events, as well as loudness fluctuation.
To enable machines to understand acoustic environments, computational analysis of audio scenes and events [5] has been studied extensively, such as the works by detecting and classifying acoustic scenes and events (DCASE) community [6]. This has led to significant advancements in the recognition of acoustic scene (AS) and audio event (AE), resulting in various methods, from machine learning methods [7] to deep learning methods, such as those based on convolutional neural networks (CNN) [8] and recurrent neural networks [9]. The depth and diversity of DCASE studies are also increasing. For example, from frame-level strong label-based to clip-level weak label-based AE detection [6][10], the resolution of AE analysis is improving; from AE recognition to audio captioning [11][12] that describes AEs in an audio clip with natural language, the analysis on AEs has been upgraded. However, numerous DCASE-related works focus on the objective attributes of sounds, such as category and temporal characteristics of AS and AE, ignoring the effects these sounds have on people, such as the different dimensions of emotion they evoke.
@@
To describe sound-related emotion perceptions, ISO/TS 12913-3:2019 [14] recommends using the soundscape circumplex model (SCM) [13], a framework inspired by the affect theory of emotions [15]. As shown in Fig. 1, the SCM is scored on eight 5-point Likert scales (pleasant, vibrant, eventful, chaotic, annoying, monotonous, uneventful, and calm) to describe the perceptual attributes of soundscapes. Some prior studies [16][17] explore the relationships between AEs and annoyance, which is one of the 8 attributes of perceived affective quality (PAQ). However, there is still a research gap between various AEs and the 8-dimensional (8D) affective qualities (AQs) in PAQ, and a larger research gap between the ASs, AEs, and affective responses to 8D AQs in PAQ.
In addition to the works above that analyse AS and AE from the perspective of classification tasks, combining audio processing with natural language processing (NLP) has recently become a research hotspot. In the audio caption task in DCASE 2020 [12], the AEs and AS in an audio clip are described with texts to convert audio content to captions. However, audio captioning focuses on the specific, objective AE or AS-related information, such as “birds are of chirping the chirping and various chirping[12]. Current audio captioning systems have not explored the listener’s response to the audio along the affective dimension, i.e., whether hearing birds chirping brings pleasure or annoyance to the listener. Despite the great progress in audio captioning for the acoustic environment of soundscapes, little attention has been paid to the affective information carried by sounds in soundscapes.
@@
To fill this research gap, we propose the soundscape captioning task where the content of audio recordings from a soundscape is described using context-aware text with perspectives of the AS, AE, and emotion-related AQ. This enables affective information exploration of anthropocentric soundscapes, as defined in ISO 12913-1:2014. Inspired by the excellent performance of large language models (LLMs) [18] on NLP tasks, we propose a LLM-based automatic soundscape captioner (SoundSCaper) system for the soundscape captioning task by integrating coarse-grained ASs, fine-grained AEs, and human-perceived AQs information. SoundSCaper integrates rich prior knowledge in the general LLM represented by the generative pre-trained transformer (GPT) [19], automatically generating captions to describe soundscape content from the perspectives of AS, AE, and AQ. In this way, SoundSCaper, which relies on the rich knowledge of LLM, can include contextual and auditory scene-based expectations when describing soundscapes, which is a crucial aspect of soundscape [4].
Specifically, to feed the LLM with precise information extracted from recordings, we propose a sound affective quality network (SoundAQnet) to simultaneously model AS and AE in acoustic environments and their corresponding AQ. A prior study explored the possibility of classifying sounds from AQ perspectives and using acoustic features derived from psycho-acoustics or music theory [20]. Although such an approach has potential benefits, associating meaning to sounds, i.e., recognizing audio events (AEs), is a more plausible pathway in human soundscape assessment. Hence, SoundAQnet adopts Mel features [21] commonly used in AE and AS recognition tasks with a single psycho-acoustic feature (instantaneous loudness defined in ISO 532-1 [22]) to capture AS, AE, and human-perceived AQ in soundscapes. We then integrate the acoustic information and corresponding AQ captured by SoundAQnet with a generic LLM as the proposed SoundSCaper system, which generates common-language captions to parse soundscapes’ semantic and affective context, rather than using only a single numerical metric to describe the soundscape as in previous soundscape works. Thus, the SoundSCaper system dedicated to the soundscape captioning task will bridge the gap between single numerical indicators and human perception, making it easier for humans to understand the soundscape’s acoustic content and affective information. To the best of our knowledge, this is the first system that simultaneously characterizes AS, AE, and emotion-related AQ in acoustic environments, and the first system that automatically generates captions describing soundscapes by combining the rich prior knowledge contained in LLM with three-view information (AS, AE, and AQ) related to sound.
Most soundscape studies rely on human listening tests and questionnaires, which are time-consuming and laborious. The successful application of soundscape captioning will promote automatic soundscape analysis, help people understand acoustic environments more deeply, create immersive virtual environments [23], and improve urban soundscape planning [24]. It can also enhance the environmental awareness of visually and hearing-impaired people, allowing them to understand and respond to changes in surroundings more easily [25].
@@
This paper strives to advance machine listening by linking it with affective computing and contextual interpretation, thus going beyond conventional recognition and classification of sounds. Our work offers the potential to enable machines to have a comprehensive and emotionally attuned perception of auditory scenes and events. The contributions of this paper are as follows: 1) We propose the soundscape captioning task, where a soundscape is described in free texts from the perspectives of AS, AE, and AQ, thus bridging the gap between audio captions and the human-perceived AQs of sounds; 2) We propose a sound affective quality network (SoundAQnet) to simultaneously model the coarse-grained AS and fine-grained AE, as well as human-perceived AQ; 3) We utilize the rich knowledge embedded in LLM about expected sounds in various scenes to develop the automatic soundscape captioner (SoundSCaper) system, which translates SoundAQnet’s predictions into clear, readable captions; thus, soundscape descriptions are no longer limited to single numerical features but extended to free texts easy to comprehend by humans; 4) To measure the quality of soundscape captions generated by SoundSCaper, we introduce the Transparent Human Benchmark for Soundscapes (THumBS) as a metric, and evaluate the performance of SoundSCaper on the test set and the mixed external dataset. We have released the code and models, human assessment data, and expert evaluation statistics to the homepage (https://github.com/Yuanbo2020/SoundSCaper).
The remaining sections are organized as follows. Section II introduces the proposed soundscape captioning task. Section III proposes the SoundSCaper system, based on SoundAQnet and LLM. Section IV presents the dataset and experimental setup of SoundAQnet, and analyses SoundAQnet’s performance. Section V discusses the internal and external datasets used for human assessment, analyzes the results of jury’s scores for soundscape captions generated by the proposed SoundSCaper system and human experts, and discusses their characteristics and differences. Section VI concludes.
II Soundscape captioning
Soundscape captioning describes the holistic experience and understanding of an acoustic environment. Identifying and recognizing AEs and assigning meaning to them is one of the cornerstones of this personal and individual experience. Because of this, soundscape evaluation questionnaires (e.g., ISO/TS 12913-2 [26]) inquire about the classes of sounds people hear. Thus, captioning should include a description of the categories of audible sounds. In addition, the holistic evaluation is also affected by psycho-acoustics-related loudness and the context of sounds. Therefore, from the perspective of holistic experience and understanding, soundscape captioning should consider the categories of sounds, the context in which they occur, and the affective quality they may evoke in humans. Among DCASE-related tasks, audio event classification (AEC) aims to tag the types of AEs of audio clips with semantic labels; acoustic scene classification (ASC) aims to identify the category of the environment in which the sound was recorded, i.e., its context. Both AEC and ASC tasks contribute to the accuracy of soundscape captioning.
Audio captioning converts audio content, mainly AEs, into text. Soundscape captioning starts from the perspective of soundscape with human perception rather than just AEs, adds AS information, and focuses on exploring emotional impact. As shown in Fig. 2, soundscape captioning is a task that aims to generate natural language sentences to describe the soundscape content, in particular, AS, AE, and human-perceived AQ. Soundscape captioning enables a machine to understand what AS the recording comes from, what AEs occur, its overall AQs, and possible emotional impact on people, and then summarizes these in human-understandable languages.
@@ II Soundscape captioning
Figure 2: Results of tasks on the same audio clip: coloured texts show differences between audio captioning and soundscape captioning. Blue texts indicate human-perceived AQ-related descriptions unique to soundscape captioning.
@@
The soundscape captioning task aims to generate a language description $\boldsymbol{D}$ based on the input audio clip $\boldsymbol{A}$. To this end, first, we extract the AS and AE information, as well as the affective response values of eight AQs in Fig. 1, i.e. pleasant, eventful, chaotic, vibrant, uneventful, calm, annoying, and monotonous [27] from the audio clip $\boldsymbol{A}$, by building an acoustic model ($am(\cdot)$), i.e., $\{AS,AE,AQ\}=am(\boldsymbol{A})$. Then, we form a textual description of the soundscape by a language model ($lm(\cdot)$), i.e., $\boldsymbol{D}=lm(AS,AE,AQ)$.
III Automatic soundscape captioner system
The proposed automatic soundscape captioner system, SoundSCaper in Fig. 3, consists of two parts: the acoustic model ($am(\cdot)$) SoundAQnet, and the language model ($lm(\cdot)$). With SoundAQnet, we can extract information about AS, AE, and the perceived AQ values, from variable-length audio recordings. With the general LLM like GPT [18], we can generate soundscape descriptions by embedding AS and AE information, and human-perceived AQ, into the text prompt.
@@ III Automatic soundscape captioner system
III-A Acoustic model: The proposed SoundAQnet
The acoustic model aims to simultaneously model the ASs and AEs in acoustic environments, as well as the corresponding affective responses to AQs, i.e., PAQ 8 attributes in Fig. 1. Since real-life audio clips are of variable length, enabling the acoustic model to handle audio clips of different lengths is also an issue that needs to be considered. In previous work, log Mel spectrogram is a commonly used acoustic feature [28], offering excellent performance in AS and AE recognition. In soundscape studies [29], loudness from psychoacoustics, related to human perception of sound level, is a non-negligible factor. Thus, both Mel and loudness are used in SoundAQnet to capture the AS, AE, and AQ in audio recordings.
Taking a 30-second audio clip as an example, following the setting of log Mel spectrum in [21], the frame length and hop size are 32ms and 10ms, respectively, resulting in Mel features having 3000 frames. The psycho-acoustics-related loudness features extracted according to the ISO 532-1:2017 standard [22] have 15000 frames. To process these long input features with few parameters, SoundAQnet uses dilated convolution [30] to obtain a large receptive field with limited computing resources. SoundAQnet needs to consider the 8 different AQs recommended by ISO/TS 12913-3:2019 [14] in Fig. 1. It is worth noting that the generation, development, and fading of emotions is a dynamic process. Thus, SoundAQnet employs multiscale convolution blocks to extract human-perceived AQs, so that it can capture representations of different time resolutions for different AQs. SoundAQnet uses a pooling operation [21] after the last convolutional layer to receive audio input of any length. The pooling operation will mitigate the impact of various lengths on representations of different resolutions extracted by multiscale convolution blocks, unifying their dimensions for the next processing.
Figure 4: The acoustic model SoundAQnet simultaneously models acoustic scene (AS), audio event (AE), and emotion-related affective quality (AQ).
III-A1 Mel-based branch
To capture the acoustic information and the corresponding AQ information at different time scales, four sub-branches use convolutional kernels of different sizes, i.e., [(3, 3), (5, 5), (7, 7), (9, 9)], applied to input features on the (time, frequency) axis, respectively. Each sub-branch comprises three convolution blocks with 16, 32, and 64 filters. Dilated convolution [30] is used to capture multiscale contextual information by obtaining a larger receptive field size (RFS) with fewer parameters. The gridding artifacts [31] of dilated convolutions can lead to compromised information continuity and loss of local information. Thus, a hybrid dilated convolution [30] scheme is used, where each convolution block uses different dilation rates to mitigate the gridding problem. The dilation rates in the three convolutional blocks are in order [(1,1), (2,1), (3,1)], allowing the branches to comprehensively extract context from a broader and more coherent receptive field. The dilation rate only varies along the time axis, as the frequency dimension is often relatively small.
In the Mel branch, each 2D convolution (Conv2D) block refers to the design of VGG [32] and consists of two convolution layers with batch normalization and ReLU function [21]. Taking the largest kernel (9, 9) in the Mel branch as an example, there are three Conv2D blocks, i.e., six 2D convolution layers, according to the convolution RFS calculation formula,

### v3

@@ Abstract
We live in a rich and varied acoustic world, which is experienced by individuals or communities as a soundscape. Computational auditory scene analysis, disentangling acoustic scenes by detecting and classifying events, focuses on objective attributes of sounds, such as their category and temporal characteristics, ignoring their effects on people, such as the emotions they evoke within a context. To fill this gap, we propose the affective soundscape captioning (ASSC) task, which enables automated soundscape analysis, thus avoiding labour-intensive subjective ratings and surveys in conventional methods. With soundscape captioning, context-aware descriptions are generated for soundscape by capturing the acoustic scenes (ASs), audio events (AEs) information, and the corresponding human affective qualities (AQs). To this end, we propose an automatic soundscape captioner (SoundSCaper) system composed of an acoustic model, i.e. SoundAQnet, and a large language model (LLM). SoundAQnet simultaneously models multi-scale information about ASs, AEs, and perceived AQs, while the LLM describes the soundscape with captions by parsing the information captured with SoundAQnet. SoundSCaper is assessed by two juries of 32 people. In expert evaluation, the average score of SoundSCaper-generated captions is slightly lower than that of two soundscape experts on the evaluation set D1 and the external mixed dataset D2, but not statistically significant. In layperson evaluation, SoundSCaper outperforms soundscape experts in several metrics on datasets D1 and D2. In addition to human evaluation, compared to other automated audio captioning (AAC) systems with and without LLM, SoundSCaper performs better on the ASSC task in several natural language processing (NLP) based metrics. Overall, SoundSCaper performs well in human subjective evaluation and various objective captioning metrics, and the generated captions are comparable to those annotated by soundscape experts. The model, source code, LLM scripts, human assessment data, instructions, and evaluation statistics are all publicly available.
@@ I Introduction
Soundscape plays a vital role in shaping our daily experience, affecting various aspects including mood, behaviours, and overall well-being [1]. Understanding how people perceive soundscape has become important in urban design, environmental psychology, and interactive technology [2]. The international standard organization (ISO) 12913-1:2014 [3] defines soundscape as: “the acoustic environment as perceived or experienced and/or understood by a person or people, in context”, which emphasizes the interaction between a person or people and the acoustic environment. Thus, a soundscape is not only about the audio events (AEs) present in the acoustic scene (AS), but also the emotions and mood evoked by the acoustic environment and events therein.
To describe the affect of soundscape, ISO/TS 12913-3:2019 [4] recommends using the soundscape circumplex model (SCM) [5], a framework inspired by the affect theory of emotions [6]. In Fig. 1, the SCM is scored on eight 5-point Likert scales to describe the perceptual attributes of soundscapes. Some prior studies [7][8] explore the relationships between AEs and annoyance, which is one of the 8 attributes of perceived affective quality (PAQ) in SCM. These perceptions are shaped by sound characteristics, contextual cues, and prior experience or common knowledge specific to culture. In addition, the affect of soundscape is highly related to the perception of AS as a whole. Therefore, to describe a soundscape, it is crucial to exploit both the physical information about the acoustic environment that can be estimated using techniques such as acoustic scene classification (ASC) and audio event classification (AEC), and the perceptual information, such as human-perceived affective quality (AQ) of the soundscape.
@@
To enable machines to understand acoustic environments, computational analysis of audio scenes and events [9] has been studied extensively, e.g., by the detection and classification of acoustic scenes and events (DCASE) community [10]. This has led to significant advancements in the recognition of ASs and AEs, resulting in various methods, from machine learning methods to deep learning methods, such as AE detection methods [10][11][12] based on frame-level strong labels or clip-level weak labels, and ASC methods [13][14][15]. More recently, natural language has been used to describe audio content, including ASs and AEs, leading to an emerging area called automated audio captioning (AAC) [16] [17].
However, the DCASE-related works focus mainly on the objective attributes of sounds, while ignoring the effects these sounds have on people, such as the different emotions they evoke. This results in the relations between various AEs and 8-dimensional (8D) affective qualities (AQs) in PAQ remaining unexplored, let alone the relations between AS, AEs, and affective responses to 8D AQs. For example, the AAC systems often describe AEs or AS-related information, such as “birds are of chirping the chirping and various chirping” in the AAC task in DCASE 2020 [17], but they have not explored the listener’s response to the audio along the affective dimension, i.e., whether hearing birds chirping brings pleasure or annoyance to the listener. Despite the substantial progress in AAC for the acoustic environment, little attention has been paid to the affective information carried by AEs in a soundscape or the soundscape as a whole. More specifically, there is a significant research gap in captioning that links the acoustic environments (ASs and AEs) and the human-perceived AQs of soundscapes.
@@
To fill this research gap, we propose a new task, called affective soundscape captioning (ASSC), where a soundscape is described using context-aware texts, detailing AS, AE, and emotion-related AQ. This enables the exploration of affective information from anthropocentric soundscapes, as defined in ISO 12913-1:2014 [3]. Inspired by recent advancements in large language models (LLMs) [18], we propose an LLM-based soundscape caption system (SoundSCaper) for the ASSC task by integrating coarse-grained ASs, fine-grained AEs, and human-perceived AQs. SoundSCaper integrates the rich prior knowledge in LLMs like generative pre-trained transformer (GPT) models [19], enabling context-sensitive and affectively meaningful descriptions of soundscapes.
To provide precise acoustic inputs to the LLM, we propose a sound affective quality network (SoundAQnet) to simultaneously model AS and AE in acoustic environments, and their corresponding AQ. SoundAQnet combines the Mel-spectrogram features [12] commonly used in ASC and AEC tasks with loudness (Zwicker loudness, defined in ISO 532-1 [20]) to capture AS, AE, and AQ in the soundscape. Loudness, which measures the subjective impression of human perception of sound, has important implications for AQ modelling, as the perceived loudness of a sound can change its contribution to affect, from gentle background to extremely disturbing sounds. Integrating the output from SoundAQnet with GPT allows SoundSCaper to generate natural language descriptions, thereby linking objective acoustic indicators (i.e., AS and AE) with human-centred perceptual indicators (i.e., PAQ). To our knowledge, we are the first to build models that simultaneously characterize AS, AE, and AQ in a soundscape and to describe the soundscape with affective captions using an LLM based on the three-view information (i.e., AS, AE, and AQ).
@@
This paper strives to advance machine listening by linking it with affective computing and contextual interpretation, thus going beyond conventional recognition and classification of sounds. Our work offers the potential to enable machines to have a comprehensive and emotionally attuned perception of auditory scenes and events. The contributions of this paper are as follows: 1) We propose the ASSC task, where a soundscape is described in free texts from the perspectives of AS, AE, and AQ, thus bridging the gap between audio captions and the human-perceived AQs of sounds. 2) We propose SoundAQnet to simultaneously model the coarse-grained AS and fine-grained AE, as well as human-perceived AQ. 3) We utilize the rich knowledge embedded in LLM about expected sounds in various scenes to develop the automatic SoundSCaper system, which translates SoundAQnet’s predictions into human-understandable natural language captions; to this end, careful soundscape-focused prompt engineering is introduced. This transforms soundscape descriptions from limited numerical values into comprehensible free text rich in acoustic context and human AQ perception. 4) To measure the quality of the soundscape captions, we introduce the Transparent Human Benchmark for Soundscapes (THumBS) as a metric and evaluate the performance of SoundSCaper on the test set and the mixed external dataset.
Next, Section II discusses related work. Section III introduces the proposed ASSC task. Section IV proposes the SoundSCaper based on SoundAQnet and LLM. Section V presents SoundAQnet experiments. Section VI presents human evaluation experiments on SoundSCaper by expert and layperson groups, compares the results of SoundSCaper and AAC systems on the ASSC task, and discusses their characteristics. Section VII concludes. We have released the code and models, human assessment data, and human evaluation statistics to the homepage (https://github.com/Yuanbo2020/SoundSCaper).
II Related Work
This section reviews related work on soundscape captioning.
II-A Audio Captioning
Audio captioning [16][17] resembles soundscape captioning when it comes to disentangling the auditory scene into separate sounds. Various automated audio captioning (AAC) systems [21, 22, 23] aim to describe AEs and physical properties of acoustic environments using text. ConvNeXt-Trans [22], the baseline of DCASE 2024 challenge [17], excels at the AAC task. The audio encoder of ConvNeXt-Trans is ConvNeXt, and its decoder consists of a Transformer decoder with a structure similar to GPT. P-LocalAFT [23] allows local information to be captured while retaining global information, which captures AEs of different durations for precise captions. These AAC systems focus on AEs and ASs, as a result, they can describe the objective information of acoustic environments. However, they fail to capture emotions and moods perceived by humans, as the datasets used for model training lack AQ labels.
In addition to AAC systems without LLM, audio-LLM-based GAMA [24] and Qwen-Audio [25] perform well on multiple audio captioning datasets. GAMA uses an acoustic model to extract the information and then feeds it to an LLM to generate captions. GAMA’s framework is similar to SoundSCaper. Although audio-LLMs [24][25] have some audio perception capability, such perception focuses on the objective content of audio and serves audio understanding and reasoning; i.e., it is not the affective perception in soundscapes, let alone the perception of 8D AQs in ISO/TS 12913-3:2019.
II-B Affective Computing
In soundscape affective computing, arousal-valence dimensional models (AVDM) [2] are often used to capture human emotional responses. However, AVDM has limited ability to distinguish different affective perceptions, making it difficult to model complex emotion states [26]. Based on the pleasantness-eventfulness framework, the performance of multiple acoustic features on urban soundscapes is analyzed, and the results [27] show that Mel-based features predict the pleasantness and eventfulness of soundscapes well. And Gammatone cepstral coefficients [28] have been shown to be feasible in assessing the valence and arousal of sounds. However, the study [28] is based on synthetic rather than real datasets. This paper models human affective responses based on a multi-user annotated real soundscape dataset according to 8D AQs in ISO standards.
II-C Soundscape Analysis
Soundscape analysis can be roughly categorized into acoustic environment-oriented [9, 10, 11, 12] and affective perception-oriented [1][5][29, 30, 31]. The former focuses on the objective acoustic environment understanding, e.g., ASC and AEC. The latter focuses on affective perceptions of soundscapes to characterize emotional states. The international affective digital sound (IADS) [29] dataset explores discrete emotional categories elicited by 167 individual stimuli, and the international soundscape database (ISD) [30] explores the relationship between annoyance and 24 categories of AEs in daily life. In addition, ARAUS [32] studies the unique subjective perceptual responses of 605 participants to 25440 augmented soundscapes presented as audiovisual stimuli.
To the best of our knowledge, there are no studies on automatic descriptions of soundscapes, especially AASC. The successful application of the proposed AASC will promote automated soundscape analysis, improve urban soundscape planning, and improve environmental awareness of visually and hearing-impaired people [33][34].
@@ II-C Soundscape Analysis
Figure 2: Results of tasks: coloured texts show differences between automated audio captioning (AAC) and affective soundscape captioning (ASSC) tasks. Blue texts indicate human-perceived AQ-related descriptions unique to ASSC.
III Affective Soundscape Captioning
This section presents problem formulation and task definition for ASSC, and the SoundScaper system for AASC.
III-A Problem Formulation and Task Definition
Affective soundscape captioning (ASSC) describes the holistic experience and understanding of an acoustic environment. Identifying and recognizing AEs and assigning meaning to them is one of the cornerstones of this personal and individual experience. Because of this, soundscape evaluation questionnaires (e.g., ISO/TS 12913-2 [35]) inquire about the classes of sounds people hear. Thus, ASSC should include a description of the relevant categories of audible sounds.
The holistic evaluation is also affected by psychoacoustics-related loudness [20] and the context of sounds. Therefore, from the perspective of holistic experience and understanding, ASSC should consider the categories of sounds, the context in which they occur, and the AQ they may evoke in humans. Fig. 2 contrasts ASSC to related tasks. The AEC task aims to tag the types of AEs of audio clips with semantic labels; the ASC task identifies the environment category where the sound is recorded, i.e., its context. The AAC task converts audio content, mainly AEs, into text. ASSC starts from the perspective of soundscape with human perception rather than just AEs, adds AS information, and focuses on emotional impact. AASC adds affective attributes to the textual description and suggests the human affect evoked by the acoustic environment.
III-B Overview of the Proposed SoundScaper System
@@
The proposed automatic soundscape captioner, SoundSCaper, as shown in Fig. 3, consists of two parts: the acoustic model ($am(\cdot)$) SoundAQnet, and the language model ($lm(\cdot)$). SoundSCaper aims to generate a language description $\boldsymbol{D}$ to describe the soundscape based on the input audio clip $\boldsymbol{A}$. To this end, first, we extract the AS and AE information, as well as the affective response values of eight AQs in Fig. 1, i.e. pleasant, eventful, chaotic, vibrant, uneventful, calm, annoying, and monotonous [32] from the audio clip $\boldsymbol{A}$, by building an acoustic model ($am(\cdot)$), i.e., $\{AS,AE,AQ\}=am(\boldsymbol{A})$. Then, we form a textual description of the soundscape by a language model like GPT [18] ($lm(\cdot)$), i.e., $\boldsymbol{D}=lm(AS,AE,AQ)$.
IV The proposed SoundSCaper system
This section introduces SoundSCaper in detail from two aspects: the acoustic model and the language model.
@@ IV The proposed SoundSCaper system
IV-A Acoustic Model: the Proposed SoundAQnet
The SoundAQnet aims to simultaneously model ASs and AEs in acoustic environments, as well as the corresponding affective responses to the soundscape, i.e., PAQ 8 attributes shown in Fig. 1. In recognition of AS and AE, the Mel spectrogram is a typical acoustic feature [12][15] with excellent performance. In soundscape studies [1][2][31], perceived loudness, approximated as calculated as Zwicker-loudness in ISO 532-1:2017 [20], is of primary importance to estimate AQ even if AEs are known. Thus, both Mel and loudness are used to capture the AS, AE, and AQ in soundscape audio clips. The novelty of SoundAQnet lies in the design of the network structure and the loss function to enable the joint modelling of ASs, AEs, and AQs, as described below.
IV-A1 Network
To process these long input features with few parameters, SoundAQnet uses dilated convolution [36] in its Mel-based and Loudness-based branches to obtain a larger receptive field size (RFS) with limited computing resources.
Figure 4: The SoundAQnet simultaneously models acoustic scene (AS), audio event (AE), and emotion-related affective quality (AQ).
1.1) Mel-based Branch: Audio events are often in different spectral-temporal scales, ranging from short and transient to long-lasting events. Therefore, the Mel-based branch employs four sub-branches to obtain multi-scale representation, by applying convolutional kernels of different sizes, i.e., [(3, 3), (5, 5), (7, 7), (9, 9)], to the input features on the (time, frequency) axis, respectively. Each sub-branch comprises three convolution blocks with 16, 32, and 64 filters. The gridding artifacts [37] of dilated convolutions can lead to compromised information continuity and loss of local information. Thus, a hybrid dilated convolution [36] scheme is used, where the dilation rates in the three convolutional blocks are in order [(1,1), (2,1), (3,1)], allowing the branches to extract context from a broader and more coherent receptive field along the time axis. The dilation rate only varies along the time axis, as the frequency dimension is often relatively small.
In the Mel branch, each 2D convolution (Conv2D) block refers to the design of VGG [38] and consists of two convolution layers. Taking the largest kernel (9, 9) as an example, there are three Conv2D blocks, i.e., six 2D convolution layers, according to the convolution RFS calculation formula,